OpenStudy (anonymous):

How do I find a horizontal tangent? (Calc 1)

OpenStudy (anonymous):

Lets say I have\[f(x)=x^3+3x^2+3x^2+3\] I know that the slope is 0, and I can derive, but how do I find y1 or x1 ?

OpenStudy (anonymous):

@eliassaab, @ganeshie8

ganeshie8 (ganeshie8):

first derivative at a point gives you the slope of tangent at that point

ganeshie8 (ganeshie8):

and you know that the slope of a horizontal tangent is 0

ganeshie8 (ganeshie8):

start by finding the derivative, set it equal to 0 and solve x

ganeshie8 (ganeshie8):

whats the derivative of f(x) ?

OpenStudy (anonymous):

\[\frac{d}{dx} (x^3+3x^2+3x+3)=3x^2+6x+3\]\[0=3x^2+6x+3=3(x^2+2x+1)=3(x+1)^2~~~~->~~~~x=-1\]

OpenStudy (anonymous):

So, since the x-coordinate is -1, then I know the y-coordinate. \[(-1)^3+3(-1)^2+3(-1)+3\]\[-1+3-3+3\]\[-1+3\]\[y_1=2\]

OpenStudy (anonymous):

\[y-y_1=0(x-x_1)\]\[y-2=0\]\[y=2~~~(answer)\]

OpenStudy (anonymous):

Yes, tnx that makes sense, https://www.desmos.com/calculator/lye8nzzwlv Appreciate it

ganeshie8 (ganeshie8):

Excellent !

OpenStudy (anonymous):

your job is indeed is excellent.

OpenStudy (anonymous):

Just what I needed.

OpenStudy (anonymous):

And thus the vertical tangent would be x=-1, (setting a vertical line through the point)

OpenStudy (anonymous):

Tnx once again:) Have a good day sir:)

ganeshie8 (ganeshie8):

no vertical tangents

OpenStudy (anonymous):

Why not?

ganeshie8 (ganeshie8):

we only have one horizontal tangent y = 2 thats all

OpenStudy (anonymous):

Ohh true dat:)

OpenStudy (anonymous):

it's +- infinity depending on whether it is going up or down.

ganeshie8 (ganeshie8):

a vertical tangent looks like this : 

OpenStudy (anonymous):

So a vertical tangent can only be, when the curve itself isn't a function?

ganeshie8 (ganeshie8):

yes the slope is infinity for a vertical tangent

ganeshie8 (ganeshie8):

thats a good question

ganeshie8 (ganeshie8):

you can get vertical tangent for functions also

ganeshie8 (ganeshie8):

let me think of one..

OpenStudy (anonymous):

not really, because that means that the slope of the function is undefined, i.e. more than 1 point on the same y value.

OpenStudy (anonymous):

(I think)

ganeshie8 (ganeshie8):